OpenStudy (anonymous):

what is the geometric mean between 25 and 144

OpenStudy (anonymous):

The geometric mean if two numbers \(a\) and \(b\) is \(\sqrt{ab}\).

OpenStudy (anonymous):

A geometric mean between 2 numbers is the square root of their product. \[\sqrt{25*144}=60\]

OpenStudy (anonymous):

of*

OpenStudy (anonymous):

thanks